### v1

# Automatic Model Extraction of the Match Standard in Symmetric–Reciprocal–Match Calibration

Abstract
This paper addresses the modeling of parasitics of the match standard in the symmetric-reciprocal-match (SRM) calibration method of vector network analyzers (VNAs). In the general SRM procedure, the match standard is assumed to be fully known. Here, we demonstrate that the match can be modeled with an arbitrary frequency-dependent model using a non-linear global optimization procedure. To highlight the validity of the suggested approach, numerical tests were conducted, demonstrating the ability to recover the match standard parasitic model down to software numerical precision. Additionally, we performed microstrip line measurements to compare the SRM calibration with match modeling to the multiline thru-reflect-line (TRL) calibration one, showing that automatic model extraction can achieve accuracy similar to using a match standard defined through multiline TRL calibration.
I Introduction
The common calibration procedures for vector network analyzers (VNAs) rely on known standards. Most common procedure is the short-open-load-thru (SOLT) method [1], or with an unknown reciprocal device in the short-open-load-reciprocal (SOLR) method [2]. VNA calibration has evolved and improved in accuracy by introducing new methods that rely on partially defined standards. Such calibration techniques are summarized as self-calibration procedures [3]. Well-known and often applied representatives of self-calibration procedures are, among others, the aforementioned SOLR method, the thru-reflect-line (TRL) and the multiline TRL [4, 5, 6, 7], the line-reflect-match (LRM), the thru-match-reflect-reflect (TMRR), and the line-reflect-reflect-match (LRRM) methods [8, 9, 10, 11].
The symmetric-reciprocal-match (SRM) method was recently introduced in [12], which relies solely on the symmetry and reciprocity properties of the standards while requiring a fully defined match standard. This method offers advantages over conventional self-calibration procedures, such as LRM/TMRR/LRRM, as it eliminates the need for a defined thru/line during calibration.
This paper extends the SRM method by incorporating automatic parasitic extraction modeling of the match standard. The enhancement requires only knowledge of the match standard’s DC resistance, while its parasitic response is modeled as a frequency-dependent equivalent circuit. Unlike the LRRM method [11], which uses a simple inductance model, the SRM approach allows for more complex match standard structures through global optimization.
A significant advantage of the proposed method is that all standards are partially defined, except for the match standard, whose DC value is needed to establish the reference impedance definition of the calibrated S-parameters.
The article is organized as follows. Section II reviews the general procedure for the SRM method. Section III discusses the parasitic modeling of the match and the optimization procedure. Section IV presents numerical analysis using synthetic data and experimental measurements on printed circuit boards (PCB), followed by conclusions in Section V.
II SRM Calibration Review
The SRM calibration method relies on measuring symmetric one-port standards without predefined characteristics, combined with a transmissive device measurement. This combination enables the formulation of an eigenvalue problem that partially solves for the VNA error boxes. The complete solution is then obtained using the reciprocity property of the transmissive device and a pair of defined match standards. The method requires a minimum of three symmetric one-port standards (typically open, short, and an arbitrary load), measured both individually and in combination with the transmissive device from either port, as illustrated in Fig. 1.
Figure 1: Two-port VNA error box model illustrating the standards used in SRM calibration. All matrices are provided as T-parameters. The index $i$ indicates the measured standard, where $i=1,2,\ldots,M$, with $M\geq 3$. Note that the last standard could also be measured at port-B.
The measurement of the transmissive device can be written in T-parameters as follows:
$\boldsymbol{M}_{\mathrm{net}}=\underbrace{k_{a}k_{b}}_{k}\underbrace{\left[\begin{matrix}a_{11}&a_{12}\\ a_{21}&1\end{matrix}\right]}_{\boldsymbol{A}}\boldsymbol{T}_{\mathrm{net}}\underbrace{\left[\begin{matrix}b_{11}&b_{12}\\ b_{21}&1\end{matrix}\right]}_{\boldsymbol{B}},$ (1)
where $\boldsymbol{M}_{\mathrm{net}}$ and $\boldsymbol{T}_{\mathrm{net}}$ represent the measured and actual T-parameters of the network, respectively. The matrices $\boldsymbol{A}$ and $\boldsymbol{B}$ represent the error boxes of the left and right ports, respectively, containing the first six error terms. The parameter $k$ identifies the seventh error term, which describes the transmission error between the ports.
The measurements of the one-port standards can be represented using Möbius transformation notation, as described below [13, 12]:
$\mathcal{M}_{\boldsymbol{A}}\left({\rho^{(i)}}\right)$ $=\Gamma_{a}^{(i)}=\frac{a_{11}\rho^{(i)}+a_{12}}{a_{21}\rho^{(i)}+1},$ (2a)
$\mathcal{M}_{\boldsymbol{P}\boldsymbol{B}^{-1}\boldsymbol{P}}\left({\rho^{(i)}}\right)$ $=\Gamma_{b}^{(i)}=\frac{b_{11}\rho^{(i)}-b_{21}}{1-b_{12}\rho^{(i)}},$ (2b)
where $\Gamma_{a}^{(i)}$ and $\Gamma_{b}^{(i)}$ are the $i$th measured reflection coefficients from the left and right ports, respectively. The actual response of the standard is denoted by $\rho^{(i)}$. The matrix $\boldsymbol{P}$ is a $2\times 2$ permutation matrix,
$\boldsymbol{P}=\boldsymbol{P}^{T}=\boldsymbol{P}^{-1}=\begin{bmatrix}0&1\\ 1&0\end{bmatrix}$. (3)
A Möbius transformation of the form $\mathcal{M}_{\boldsymbol{Q}}\left({z}\right)$ takes the elements of a $2\times 2$ matrix $\boldsymbol{Q}$ and maps the input variable $z$ through a linear fractional transformation. This is written in general as follows:
$\mathcal{M}_{\boldsymbol{Q}}\left({z}\right)=\frac{q_{11}z+q_{12}}{q_{21}z+q_{22}};\qquad\boldsymbol{Q}=\begin{bmatrix}q_{11}&q_{12}\\ q_{21}&q_{22}\end{bmatrix}$ (4)
The reason for expressing the one-port measurements as Möbius transformations is that the operations can be composed, which translates to a simple matrix product for the coefficients [14]. This allows us to easily express the inverse of the operation by taking the inverse of the matrix of the coefficients. For example, for the port-B measurement from (2b), the input coefficient can be reverse calculated as follows:
$\mathcal{M}_{\boldsymbol{P}\boldsymbol{B}\boldsymbol{P}}\left({\Gamma_{b}^{(i)}}\right)=\frac{\Gamma_{b}^{(i)}+b_{21}}{b_{12}\Gamma_{b}^{(i)}+b_{11}}=\rho^{(i)}.$ (5)
Further details on Möbius transformation and its properties can be found in [14]. What is important here is that the measurement from port-A can be combined with the measurement from port-B, as they measure the same standard $\rho^{(i)}$ due to the imposed symmetry property, as follows:
$\Gamma_{a}^{(i)}=\mathcal{M}_{\boldsymbol{A}}\left({\mathcal{M}_{\boldsymbol{P}\boldsymbol{B}\boldsymbol{P}}\left({\Gamma_{b}^{(i)}}\right)}\right)=\mathcal{M}_{\underbrace{\boldsymbol{A}\boldsymbol{P}\boldsymbol{B}\boldsymbol{P}}_{\boldsymbol{H}}}\left({\Gamma_{b}^{(i)}}\right)$ (6)
Therefore, with at least three symmetric one-port measurements, it is possible to retrieve the coefficients of the matrix product $\boldsymbol{H}=\boldsymbol{A}\boldsymbol{P}\boldsymbol{B}\boldsymbol{P}$ through a least-squares or nullspace formulation [12].
Similarly to (6), another Möbius transformation composition can be written using the one-port measurement cascaded with the transmission network at port-A with the measurement at port-B as follows:
$\Gamma_{a,n}^{(i)}=\mathcal{M}_{\underbrace{\boldsymbol{A}\boldsymbol{T}_{\mathrm{net}}\boldsymbol{P}\boldsymbol{B}\boldsymbol{P}}_{\boldsymbol{F}}}\left({\Gamma_{b}^{(i)}}\right),$ (7)
where $\boldsymbol{F}$ is also solved using a least-squares or nullspace solution.
Therefore, with $\boldsymbol{H}$ and $\boldsymbol{F}$ and the measurement of the unknown transmissive device $\boldsymbol{M}_{\mathrm{net}}$ from (1), it can be shown that a “virtual” thru standard can be created that can be used to formulate an eigenvalue problem for both port-A and port-B coefficients. Reference [12] discusses this in greater detail, and provides the equations for the case of measuring the cascaded one-port standards at port-B.
In general, the structure of the obtained eigenvalue problem at each port is given by the following expressions:
$\frac{k}{\nu}\boldsymbol{A}\boldsymbol{P}\boldsymbol{A}^{-1}$ $=\boldsymbol{W}_{a}\boldsymbol{\Lambda}\boldsymbol{W}_{a}^{-1},\quad\forall k,\nu\neq 0$ (8a)
$\frac{k}{\nu}\boldsymbol{B}^{T}\boldsymbol{P}\boldsymbol{B}^{-T}$ $=\boldsymbol{W}_{b}\boldsymbol{\Lambda}\boldsymbol{W}_{b}^{-1},\quad\forall k,\nu\neq 0$ (8b)
where $\boldsymbol{W}_{a}$ and $\boldsymbol{W}_{b}$ are the eigenvectors to be solved for, and $\boldsymbol{\Lambda}$ contain the eigenvalues.
The last step is to incorporate the defined match standard to formulate a nullspace problem, as shown below:
$\left[\begin{matrix}-1&-1&w_{11}^{(a)}&w_{11}^{(a)}\\ 1&-1&-w_{12}^{(a)}&w_{12}^{(a)}\\ -\rho_{a}^{(m)}&-1&\Gamma_{a}^{(m)}\rho_{a}^{(m)}&\Gamma_{a}^{(m)}\end{matrix}\right]\left[\begin{matrix}a_{11}\\ a_{12}\\ a_{21}\\ 1\end{matrix}\right]=\boldsymbol{0}$ (9)
The system of equations for port-B can be obtained in a similar way, resulting in the following system of equations:
$\left[\begin{matrix}-1&-1&w_{11}^{(b)}&w_{11}^{(b)}\\ 1&-1&-w_{12}^{(b)}&w_{12}^{(b)}\\ -\rho_{b}^{(m)}&1&-\Gamma_{b}^{(m)}\rho_{b}^{(m)}&\Gamma_{b}^{(m)}\end{matrix}\right]\left[\begin{matrix}b_{11}\\ b_{21}\\ b_{12}\\ 1\end{matrix}\right]=\boldsymbol{0}$ (10)
where $\rho_{a}^{(m)}$ and $\rho_{b}^{(m)}$ are the defined match standards at both ports (they do not need to be equal), and $\Gamma_{a}^{(m)}$ and $\Gamma_{b}^{(m)}$ are their corresponding measurements, respectively. The coefficients $w_{ij}^{(a)}$ and $w_{ij}^{(b)}$ are the coefficients from the eigenvectors for both ports. Lastly, the final error term $k$ is solved using the unknown-thru procedure as in the SOLR procedure [2].
In the entire process of solving for the error terms, only the match standard needs to be explicitly defined, while all other standards are characterized through symmetry and reciprocity. The discussion in Section III addresses how to solve (9) and (10) without explicitly defining the match standard, but instead through an equivalent circuit modeling to capture its parasitic response.
III Automatic Parasitic Model Extraction
In the derivation outlined in Section II, (9) and (10) were obtained, which can only be solved if a fully defined match standard is used. In many cases, the match standard can be challenging to accurately define at higher frequencies, as parasitic elements become more impactful as the wavelength approaches the device’s physical size. This is further demonstrated with measurements in Section IV.
In general, the parasitic behavior of the match standard is more complex than just a series inductance or shunt capacitance. These non-idealities can come as a mixture of series and shunt behavior and may even include a transmission line segment. An example of such a model is depicted in Fig. 2. In this model, the DC resistance of the match standard is known, as well as the length of the offset. The remaining parameters are modeled as frequency-dependent with unknown constants. The lumped elements $L(f)$ and $C(f)$ can be modeled using polynomials, whereas the transmission line parameters, i.e., propagation constant $\gamma(f)$ and characteristic impedance $Z_{c}(f)$, can be modeled using corresponding analytical or semi-analytical equations of the physical cross-section of the transmission line.
Figure 2: Equivalent circuit model of a non-ideal match standard. The values of $l$ and $\rho_{\mathrm{dc}}$ are assumed to be known. The parameters $L(f)$, $C(f)$, $\gamma(f)$, and $Z_{c}(f)$ are modeled as frequency-dependent.
It should be noted that the DC resistance needs to be defined, as otherwise, it would not be possible to define the reference impedance. With knowledge of the DC resistance, the definition of the match standard at DC can be anchored. Also, the length of the transmission line needs to be defined for the same reason, as otherwise, a scaling ambiguity in the definition of $\gamma(f)$ would arise.
While the equivalent circuit model presented in Fig. 2 is an example of a match standard, it could also be used to describe non-ideal open or short standards, where $\rho_{\mathrm{dc}}=1$ is for an open standard and $\rho_{\mathrm{dc}}=-1$ is for a short standard. However, in general, the complexity of the model depends on how the match standard is implemented in reality. For example, in the measurement example discussed in Section IV-B, a flip-chip resistor is used, which needed to be described with a more complex lumped model to account for the mounting pads and the resistor’s internal parasitics.
The question is: if a one-port device can be described using an equivalent circuit model, how can the unknown parameters be solved without the match standard’s measurement? This can be addressed through non-linear optimization of the nullspace problem through singular value decomposition (SVD). Recalling the system of equations earlier in either (9) or (10), this system can be expanded by including other one-port standards besides the match standard, as shown in (11) below for port-A:
$\underbrace{\left[\begin{matrix}-1&-1&w_{11}^{(a)}&w_{11}^{(a)}\\ 1&-1&-w_{12}^{(a)}&w_{12}^{(a)}\\ -\rho_{a}^{(i)}(\boldsymbol{\theta}_{i})&-1&\Gamma_{a}^{(i)}\rho_{a}^{(i)}(\boldsymbol{\theta}_{i})&\Gamma_{a}^{(i)}\\ \vdots&\vdots&\vdots&\vdots\end{matrix}\right]}_{\boldsymbol{G}_{a}(\boldsymbol{\theta})}\left[\begin{matrix}a_{11}\\ a_{12}\\ a_{21}\\ 1\end{matrix}\right]=\boldsymbol{0}$ (11)
where $\Gamma_{a}^{(i)}$ and $\rho_{a}^{(i)}(\boldsymbol{\theta}_{i})$ represent the raw measurement and reflection coefficient, respectively, of the $i$th load standard. The vector $\boldsymbol{\theta}_{i}$ contains all unknown frequency-independent parameters that need to be determined, with $\boldsymbol{\theta}=[\boldsymbol{\theta}_{1}^{T},\boldsymbol{\theta}_{2}^{T},\ldots]^{T}$. It is important to note that the match standard must be part of the $i$th standards, along with at least one other one-port standard, such as a short, open, or arbitrary non-match impedance, or all of them simultaneously. Therefore, $i>1$ is required to have an overdetermined system of equations.
Since $\boldsymbol{G}_{a}(\boldsymbol{\theta})$ must have a rank of 3 to be solvable, and its nullspace corresponds to the solution vector, the optimal choice of $\boldsymbol{\theta}$ should result in a minimal fourth singular value (ideally zero), which corresponds to the nullspace [15]. Therefore, the fourth singular value corresponding to the nullspace of $\boldsymbol{G}_{a}(\boldsymbol{\theta})$ can be calculated at each frequency point and used to determine the parameter vector $\boldsymbol{\theta}$ that minimizes this value across all frequency points. This optimization procedure over the parameter vector $\boldsymbol{\theta}$ can be expressed as minimizing the average fourth singular value across all frequency points, described as follows:
$\boldsymbol{\theta}_{\mathrm{opt}}=\underset{\boldsymbol{\theta}}{\operatorname{argmin}}\;\frac{1}{N}\sum_{i=1}^{N}\sigma_{4}(\boldsymbol{\theta},f_{i})$ (12)
where $f_{i}$ refers to the $i$th frequency point, and $N$ is the total number of frequencies. Once $\boldsymbol{\theta}_{\mathrm{opt}}$ have been found, the parameters associated with the match standard are used to define its reflection coefficient accurately.
The optimization procedure in (12) can only be solved when dealing with an overdetermined system of equations, which requires having at least one additional measurement besides the match standard. This can be either a short or open standard that was assumed unknown during calibration. These standards are modeled similarly to the match standard. For the optimization itself, since singular values do not necessarily form a convex problem, a bounded global optimization method, e.g., differential evolution (DE) [16] needs to be used.
After $\boldsymbol{\theta}_{\mathrm{opt}}$ is found, the coefficients associated with the match standard are substituted back into (9) and (10). If using symmetric match pairs, optimization can be performed on only one side. However, for asymmetric match standards, each port must be solved independently. Furthermore, while additional unknown standards can be included in the SVD procedure (e.g., both short and open standards), this significantly increases the number of parameters to be solved, as the unknown parameters of each standard are not coupled. This makes it challenging even for global optimization methods to find all parameters simultaneously. Therefore, it is best practice to minimize the complexity of the optimization whenever possible, either through number of solved parameters or coupling the parameters in both ports. Additionally, the speed of the optimization is also a function of the evaluation of the models. For example, incorporating a transmission line segment, as in the numerical example in Section IV-A, can be computationally expensive when using special integral functions as the Bessel functions [17].
IV Experiments
This section discusses two experiments. The first experiment involves numerical analysis using synthetic data to demonstrate the accuracy of fitting the match standard using a nonlinear optimization procedure compared to using an ideal definition for the match standard. The second experiment involves PCB measurements to validate the automatic parasitic model extraction procedure against multiline TRL calibration.
IV-A Numerical Analysis
The numerical analysis involves creating synthetic data of co-planar waveguide (CPW) standards using the model developed in [17, 18, 19]. To accurately simulate an on-wafer setup, error boxes extracted from actual on-wafer measurements are utilized using multiline TRL calibration on a CPW impedance substrate standard (ISS). For detailed information about the measurement setup, calibration substrate, and measurements, readers should refer to [7]. The goal is to generate SRM standards based on the CPW model and embed them within the error boxes of an actual VNA setup. Fig. 3 illustrates the experimental setup.
Figure 3: Block diagram illustration of the numerical simulation concept to generate realistic synthetic data.
The CPW cross-section structure has the following dimensions: signal width of $49.1\,\mu\mathrm{m}$, ground width of $273.3\,\mu\mathrm{m}$, conductor spacing of $25.5\,\mu\mathrm{m}$, and conductor thickness of $4.9\,\mu\mathrm{m}$. The substrate is Alumina with a real-part relative permittivity of $\epsilon_{r}^{\prime}=9.9$ and loss tangent of $\tan\delta=0.0002$. The conductor is gold with relative conductivity to copper of $\sigma_{r}=0.7$, where copper conductivity is $\sigma=58\,\mathrm{MS/m}$.
The one-port SRM standards were implemented as non-ideal match, short, and open standards with a $200\,\mu\mathrm{m}$ offset, as shown in Fig. 4. For the network-load standards, a $4\,\mathrm{mm}$ line is used as the reciprocal standard, combined with the non-ideal match, short, and open standards. Note that the CPW has a complex-valued, frequency-dependent characteristic impedance that deviates slightly from $50\,\Omega$. Thus, the CPW line segments introduce both an offset and impedance mismatch, as $Z_{\mathrm{ref}}=50\,\Omega$ is used as the reference impedance for the lumped elements. The device under test (DUT) used for verification is a stepped impedance using the same CPW structure but with a $15\,\mu\mathrm{m}$ signal width.
Figure 4: Models used to simulate non-ideal load standards (a) $50\,\Omega$ match standard with $L_{0}=25\,\mathrm{pH},C_{0}=1\,\mathrm{fF}$, (b) short standard with $L_{0}=30\,\mathrm{pH},L_{1}=10\times 10^{-12}\,\mathrm{pH/Hz},C_{0}=0.5\,\mathrm{fF}$, and open standard with $C_{0}=15\,\mathrm{fF},C_{1}=5\times 10^{-15}\,\mathrm{fF/Hz},L_{0}=1\,\mathrm{pH}$. All standards include a $200\,\mu\mathrm{m}$ CPW line offset. These models are used for automatic parasitic extraction.
Only the match and short standards have been used for automatic model parameters estimation. All lumped elements and material properties ($\epsilon_{r}^{\prime}$, $\tan\delta$, and $\sigma_{r}$) were assumed unknown. While including the open standard would have been possible, it would have increased the parameter space from 11 parameters (five from the match standard and six from the short standard) to 17 parameters, significantly increasing optimization time. Symmetric models at both ports were assumed to improve optimization efficiency, as asymmetric models would double the parameter count. This consideration is particularly important since the CPW model, though analytical, uses computationally intensive Bessel functions. In the measurement example in Section IV-B, devices have been modeled using only lumped elements without offset, allowing for more parameters and independent port solutions due to faster evaluation.
Though the CPW material parameters are common to both standards, they have been treated as independent during optimization to demonstrate the procedure’s accuracy in parameter recovery.
To validate calibration accuracy, a relative error metric has been defined:
$\text{Relative Error}(x)=\left|\frac{x_{\mathrm{est}}-x_{\mathrm{true}}}{x_{\mathrm{true}}}\right|$ (13)
where $x_{\mathrm{est}}$ represents the estimated quantity and $x_{\mathrm{true}}$ is the true value.
Data processing was conducted using Python with the scikit-rf [20] and scipy [21] packages, employing the DE optimization procedure [21]. Fig. 5 shows the calibrated DUT results comparing two cases: using an ideal match standard versus using nonlinear optimization to fit unknown parameters. This ideal match models the match response by a frequency-independent $50\,\Omega$ resistor, even if the match for the simulation is calculated based on the model introduced in Fig. 4, a). The assumption of frequency independence of the match is well applicable for frequencies below $30\,\mathrm{GHz}$. Above this frequency, the ideal match standard definition produces significant errors in both $S_{11}$ and $S_{21}$. The optimization approach achieves much better accuracy, limited only by numerical precision. Fig. 6 demonstrates this by showing the relative error of estimated parameters converging to floating-point precision. It should be noted that due to the computationally expensive CPW model, the 1000 iterations shown in Fig. 6 took approximately 25 minutes to complete (using 150 frequency points).
Figure 5: True S-parameters of the stepped impedance line (DUT) compared with SRM calibration results using ideal match definition versus model fitting procedure.
Figure 6: Relative errors in estimated parameters for short and match standards during DE optimization versus iteration count.
IV-B PCB Measurement
The experimental measurement compares multiline TRL calibration against SRM calibration with and without automatic model extraction. The calibration kits are fabricated on a PCB with microstrip interface. The substrate is based on ISOLA Tachyon-100G [22] with height of $100\,\mu\mathrm{m}$. The microstrip has a strip width of $230\,\mu\mathrm{m}$ and a copper thickness of $18\,\mu\mathrm{m}$. Measurements were conducted on a probe station using ground-signal-ground (GSG) probes with $150\,\mu\mathrm{m}$ pitch. The grounded CPW pads to microstrip transition design follows principles from [23].
For multiline TRL standards, line lengths of $\{0,0.5,4,5.5,6.5,8.5\}\,\mathrm{mm}$ have been used, where the first line is the thru line. The reflect standard was implemented as an open standard. For SRM standards, symmetric open, short, and match standards and network-load with an offset line segment were implemented. The match standard uses a $50\,\Omega$ flipchip Vishay CH0402-50RGFPA [24], specified up to $50\,\mathrm{GHz}$ with $\pm 2\%$ DC resistance uncertainty. Our DC measurement showed a resistance value of $49\,\Omega$. The small size of the resistor (footprint size 0402) made manual soldering challenging, introducing asymmetry from solder bumps. Fig. 7 shows the measurement system and soldered resistors.
Figure 7: Probe station measurement setup showing the PCB containing both multiline TRL and SRM calibration kits and DUT. Inset photos show microscope views of the soldered flipchip resistors.
For comparison between SRM and multiline TRL, a stepped impedance microstrip line was used as DUT with $100\,\mu\mathrm{m}$ strip width. For proper comparison, the multiline TRL calibration was renormalized to a constant reference impedance of $50\,\Omega$. Fig. 8 shows the fitting models used for short, open, and match standards in SRM calibration. Both open and short standards were included in the optimization since they only comprise lumped elements and their evaluation is computationally efficient. The fitting was performed independently at each port. The total optimization time for both ports was approximately four minutes, fitting 22 parameters (using 197 frequency points). Only the parameters associated with the match results are used in the final SRM calibration. The resistor model follows the datasheet specifications [24], with the VIA termination modeled as an inductor.
Figure 8: Equivalent circuit model for (a) short, (b) open, and (c) match standards on the PCB. No offset is considered as the calibration plane is located at the device.
Fig. 9 compares the S-parameters of the calibrated DUT between multiline TRL and SRM. The DUT response from multiline TRL calibration was used as the reference for calculating relative error based on (13). For SRM calibration, several scenarios were examined for defining the match standard: First, assuming ideal zero reflection (typical choice when frequency-dependent response is unknown); second, using our optimization procedure from Section III to fit the models in Fig. 8; and finally, using the measurement of the match standard based on the multiline TRL calibration.
Results in Fig. 9 demonstrate that assuming an ideal match response leads to substantial errors, particularly above $10\,\mathrm{GHz}$ in $S_{21}$ magnitude. In contrast, the optimization approach yields errors comparable to defining the match standard based on multiline TRL. Some discrepancies exist between the DUT response based on SRM and multiline TRL calibrations, especially near $50\,\mathrm{GHz}$. This deviation may arise because the match standard exhibits behavior similar to a short standard near $50\,\mathrm{GHz}$, as evident in Fig. 10, thereby increasing the sensitivity of SRM calibration due to insufficient unique standards. Additionally, the discrepancy could stem from inconsistent resistor soldering, as asymmetry in the standards affects the calibration as investigated in [12]—an error that becomes amplified when the match standard behaves like a short standard near $50\,\mathrm{GHz}$. Nevertheless, achieving results comparable to multiline TRL using manually soldered flip-chip resistors is significant, considering only the DC resistance was specified during the SRM calibration. A better agreement with multiline TRL could be achieved using machine-placed or ultrasonic ball-bonded resistors, where pad parasitics are less influential.
Figure 9: S-parameters comparison of the calibrated DUT, where the relative error is computed relative to the DUT calibrated with multiline TRL.
The reflection coefficient of the match standard is presented in Fig. 10, demonstrating strong agreement between the SRM-optimized model and the multiline TRL measurement. A significant advantage of the SRM method is that it requires only the DC resistance value as input, with all other parameters determined through optimization. This approach provides greater flexibility compared to LRRM calibration, as it allows arbitrary complexity in the models for match, short, and open standards, whereas LRRM imposes restrictions on the open and match standard models [11, 25].
Figure 10: Calibrated reflection coefficient of the match standards comparing the SRM fit vs TRL measurement. Estimated parameters for the match standard model are listed in Table I.
Tables I and II list the extracted model parameters for the match standard, open, and short standards from the SRM procedure for both ports. The similar values between port-A and port-B are due to the fact that these standards were designed to be symmetrical.
TABLE I: SRM estimated parameters for the match standard model based on schematic in Fig. 8c.
$C_{1}$
$L_{1}$
$C_{2}$
$L_{2}$
$C_{R}$
$L_{R}$
$L_{via}$
     Port-A 57.37 37.80 1.50 46.69 102.02 223.01 6.54
     Port-B 53.68 50.60 1.59 28.95 108.55 216.46 4.52
TABLE II: SRM estimated parameters for the short and open models based on schematic in Fig. 8a and 8b.
Short model Open model
$C_{short}$
$L_{short}$
$C_{open}$
$L_{open}$
     Port-A 281.54 6.57 8.34 0.0095
     Port-B 199.67 6.48 8.73 0.0068
V Conclusion
In this paper, the capabilities of the SRM calibration procedure were extended to enable automatic frequency-dependent parasitic model extraction of the match standard during calibration. The model complexity can be arbitrarily defined by the user, provided that the DC value of the match is known and the number of unknown parameters is less than the number of frequency points. A global nonlinear optimization procedure finds the parameters that minimize the smallest singular value of the nullspace problem.
The capability of the SRM method was investigated to recover unknown parameters of the match standard through numerical simulation. The results demonstrated the ability to retrieve multiple parameters, including polynomial-based lumped elements and material properties of transmission line segments. Furthermore, the proposed approach was validated through measurements on a PCB using flip-chip resistors, which were modeled to account for both the non-ideality of the resistors themselves and the mounting pads, as well as the termination via arrays. The extracted model yielded results comparable to those obtained using a defined match standard measured with multiline TRL calibration.
The advantage of our proposed procedure over other calibration procedures that partially define the match standard, such as the LRRM method, lies in its modeling flexibility. While the LRRM method models the match using only an inductor, the SRM procedure can implement any combination of lumped elements or transmission lines, provided that the number of unknown parameters remains less than the frequency points. Furthermore, the SRM method offers additional advantages as all standards are defined through symmetry and reciprocity. With automatic model extraction of the match standard, only the DC resistance needs to be defined during calibration. However, one limitation of the SRM method compared to LRRM or LRM methods is the requirement for more partially defined standards, necessitating additional measurements.
Acknowledgment
The authors thank AT&S, Leoben, Austria, for the production of the PCBs used in this paper. The financial support by the Austrian Federal Ministry for Digital and Economic Affairs and the National Foundation for Research, Technology, and Development is gratefully acknowledged.